# Dual-component stellar assembly histories in local elliptical galaxies via MUSE

### v1

@@
Elliptical galaxies often exhibit complex assembly histories, and are presumed to typically form through a combination of rapid, early star formation and subsequent accretion of material, often resulting from mergers with other galaxies. To investigate theories of spheroidal galaxy formation, the objective of this work is to analyse the star formation histories (SFHs) of a sample of three isolated elliptical galaxies in the local Universe observed with MUSE at . With BUDDI, we decompose the integral field unit (IFU) datacubes into two components with Sérsic profiles, which roughly correspond to the two phases of in-situ and ex-situ star formation. To constrain the mode of growth in these galaxies, we derived the mass and light-weighted stellar ages and metallicities, and created the 2D stellar population maps of each component using pPXF. We reconstructed the mass and light-weighted SFHs to constrain the contribution of different stellar populations to the mass and luminosity of the components through cosmic time. Our results show that the ellipticals in this sample have experienced an early and rapid phase of star formation either through a rapid dissipative collapse or gas-rich major mergers concentrated in the inner component, which contributes to % of the galaxy stellar mass. The co-dominant outer component, however, had assembled the bulk of its stellar mass shortly after the inner component did, through accretion via dry mergers and possible gas accretion. This premise is supported by our observations of the inner component being primarily composed of old and metal-rich stars. The outer component has a combination of old and intermediate-age stars, with a moderate spread in metallicities. These results are analysed through the lens of the two-phase scenario, a framework developed over the years to explain the formation histories of elliptical galaxies.
@@
Elliptical galaxies are a part of the early-type galaxies (ETG) that are characterised by their spheroidal appearance and seemingly featureless and plain morphology, lacking any intricate structure. In the local Universe, they are thought to be simple and massive entities that are “red and dead”, hosting old and quiescent stellar populations and meagre amounts of gas to support star formation. However, in reality, this morphology encompasses a wide range of stellar masses and physical sizes, from compact and dwarf ellipticals to the most massive giant ellipticals. Each type has distinct structural properties and surface brightness profiles, providing valuable insights into galaxy formation and evolution. The formation scenarios for elliptical galaxies highlight the complex interplay of physical processes, from rapid monolithic collapse to gradual hierarchical merging and accretion. For instance, an often-cited mechanism for the formation of massive ellipticals is gas-rich mergers between spiral galaxies (Toomre, 1977; Hopkins et al., 2008). These processes essentially shape the structure of the elliptical galaxies, and result in different stellar populations, stellar orbits, and kinematics. Modelling the surface brightness profiles of ellipticals is therefore a crucial step in uncovering their origins and evolutionary history (MacArthur et al., 2008; Vulcani et al., 2014; Spavone et al., 2017, 2021).
@@
Elliptical galaxies were traditionally modelled with a de Vaucouleurs profile (de Vaucouleurs, 1948), which is an law. This implies a galaxy with a central concentration which becomes more extended radially outward. While this captures the light profile of many ellipticals, improved instrumentation and spatial resolution over the years have revealed a large diversity of elliptical galaxies deviating from this law (Kormendy, 1977; King, 1978). The Sérsic profile (an law) was later introduced (Sersic, 1968), which modelled the surface brightness of the ellipticals more effectively, especially the inner regions of the galaxies. This was possible due to the freedom allowed for the Sérsic index , which reverts to the de Vaucouleurs profile at . This flexibility that is incorporated into the Sérsic law is crucial because although ellipticals are indeed known for their high central concentration, the degree of concentration can vary from galaxy to galaxy. However, even the Sérsic profile might not fully represent the complex substructure that ellipticals are now expected to exhibit (Huang et al., 2013a, b; Lacerna et al., 2016), challenging the simple structure often assumed.
@@
Deep imaging has allowed the observation of internal sub-structures in elliptical galaxies, including low surface brightness tidal features (van Dokkum, 2005; Tal et al., 2009; Lazar et al., 2023) stemming from their outskirts in cluster environments, or shells (Malin & Carter, 1980; Athanassoula & Bosma, 1985; Colbert et al., 2001; Bílek et al., 2016) around the galaxies. Substructures in the central regions of ellipticals have also been observed at different redshifts, such as embedded nuclear discs (Graham et al., 2016; Dullo et al., 2018; Lacerna et al., 2020; Deeley et al., 2023). Furthermore, ellipticals have been systematically observed to fall under one of two classes based on the steepness of their central surface brightness profiles: “core ellipticals” with nearly flat central light profiles and “cuspy ellipticals” with sharply peaked central light profiles. The core and cusp regions are considered to be photometrically distinct components of ellipticals (Huang et al., 2013a). This central component therefore deviates from the standard Sérsic law that attempts to describe the surface brightness profile of the entire elliptical galaxy with a single model, without accounting for the inner component. It is therefore important to decompose the components of deceptively simple elliptical galaxies to discern their physical relevance, their formation histories, and stellar populations. For instance, Huang et al. (2013a) found that their sample of 94 nearby elliptical galaxies in the Carnegie-Irvine Galaxy Survey (CGS) required three or even four components to optimise the modelling of the surface brightness profile. Apart from the structural analysis, a complementary study in Huang et al. (2013b) revealed the underlying physical implications of modelling multiple components. Similarly Lacerna et al. (2016) found that their sample of 89 local elliptical galaxies at were best modelled with two to four components. More recently, Lima Neto et al. (2020) demonstrated that two or three components were required to model the ellipticals in the NGC 4104 fossil group. The mechanisms leading to the formation of each component are more complex than can be explained by one of the two simple scenarios involving dissipational processes as previously mentioned.
@@
The two major pillars of elliptical galaxy formation scenarios are the rapid dissipative collapse and hierarchical clustering. Over the decades, both numerical simulations and observations have been able to satisfactorily explain the structure of elliptical galaxies based on these theories. In the monolithic collapse scenario (Larson, 1974; Chiosi & Carraro, 2002), an elliptical galaxy is formed by the rapid gravitational collapse of a primordial gas cloud, followed by a substantial dissipation of energy. This theory predicts that ellipticals emerge at high redshifts () directly from the gas, with in-situ stellar mass build-up, rather than from accretion of stars formed ex-situ (Ogando et al., 2005). The rapid monolithic collapse scenario has been corroborated by strong negative metallicity gradients observed in elliptical galaxies (Domínguez Sánchez et al., 2019; Parikh et al., 2019; Goddard et al., 2017; Riffel et al., 2023), owing to the fact that the metals created during supernova explosions are retained by a strong enough gravitational potential that pushes the metal-enhanced gas toward the centre of the galaxy where the new stars form at a higher metallicity compared to the outer regions of the galaxy. The hierarchical clustering scenario (Toomre & Toomre, 1972; Cole et al., 2000; Hopkins, 2009; Avila-Reese et al., 2014) on the other hand, predicts that the present-day galaxies formed through major and minor mergers and accretion of galaxies over time. This theory has also been supported by simulations and observations where a flat or shallow metallicity gradient was exhibited by ellipticals (González Delgado et al., 2015; Taylor & Kobayashi, 2017; Benedetti et al., 2023), instead of the steep negative gradient expected from the monolithic collapse. Pure merger events are presumed to obliterate traces of the previous stellar history of the galaxy, including metallicity gradients.
@@
To better explain the mixed outcomes of elliptical galaxy observations, a hybrid scenario was proposed (Kormendy, 1989), where both mechanisms play a role in the formation of these remarkably complex galaxies. In the last decades, a two-phase scenario has been gaining traction, challenging the classical pictures of elliptical galaxy formation from either mergers or dissipative gas collapse (Kormendy, 1989). The two-phase scenario (Oser et al., 2010; Johansson et al., 2012; Huang et al., 2013a, b) suggests that the first phase of dissipational processes dominate at high redshift through cold accretion and wet mergers of gas-rich galaxies, and later at lower redshifts, the galactic evolution is driven by delayed non-dissipational dry mergers. The presence of “red nuggets” at high redshifts detected in several studies (Cimatti et al., 2004; Daddi et al., 2005; Trujillo et al., 2007; Toft et al., 2007; van Dokkum et al., 2008; Damjanov et al., 2009) already indicate an early onset of rapid star formation that settles into a compact core. Moreover, a smaller population of red nuggets that have survived to the present day has been identified, which have evolved in isolation without perturbations or interactions (Ferré-Mateu et al., 2017; D’Ago et al., 2023; Siudek et al., 2023; Moura et al., 2024). Most of these red nuggets eventually accumulate more stellar mass by accreting and merging with smaller systems to acquire the structure that is observed in present-day elliptical galaxies (Miller & Smith, 1980; Naab et al., 2006; Oser et al., 2010, 2012), while some ellipticals seem to have preserved their compact high- progenitor (Barbosa et al., 2021).
@@
Given the intricate structure of elliptical galaxies as both observations and numerical simulations clearly support, we delve into the multiple-component context of these galaxies in this paper. We make use of integral field spectroscopy (IFS) with the Multi Unit Spectroscopic Explorer (MUSE) observations of three elliptical galaxies in the local Universe, and perform 2D spectro-photometric decomposition to retrieve their structural parameters, as well as to extract their spectra and stellar populations to interpret their physical significance. The decomposition of a sample of three galaxies serves as a pilot study for further exploring the diverse properties that are inherent to elliptical galaxies over a range of stellar masses, redshifts, and environments.
@@
This paper is organised as follows. Sect. 2 outlines the data and instrument used in the IFS observations for this study. Sect. 3 describes the methods of galaxy decomposition with BUDDI, full spectral fitting with pPXF, and Voronoi binning with Vorbin. The results of this work are highlighted in Sect. 4 in the context of structural parameters, the stellar populations analysis, and star formation histories of each component. Sect. 5 provides a discussion on the physical implications and future directions of this study, and we draw our conclusions in Sect. 6. Throughout this work, we adopt a flat CDM cosmology with (Planck Collaboration et al., 2016).
@@ 2 Data and observations
MUSE (Bacon et al., 2010) is an optical integral-field spectrograph mounted at the 8.2m ESO Very Large Telescope (VLT) in Chile. It covers the nominal wavelength range of Å with spectral sampling of 1.25 Å and a field of view, and spatial resolution of 0.2″/pixel in the Wide Field Mode (WFM). The MUSE spectrograph has a spectral resolving power ranging from R1770 at 4800 Å to R3590 at 9300 Å. The three ellipticals were observed under Program ID 099.B-0411 (PI. Johnston) between July and October 2017 using the wide field mode with no adaptive optics (WFM-NoAO-N). The targets initially proposed for observation consisted of 48 galaxies selected from SDSS DR7, representing all major Hubble types from Nair & Abraham (2010). The selected targets spanned a range of Sérsic indices, B/T ratios, with an inclination as derived from Simard et al. (2011), which allowed to clearly discern the components of the galaxies. Moreover, the sample was chosen to only include unbarred galaxies, with no signatures of recent interactions, such that each galaxy can be cleanly decomposed into its major components. However, being a filler programme observed under sub-optimal weather conditions, only a small set of five elliptical galaxies with a higher set priority were observed. From this, we consider only three galaxies suitable for analysis in this work, as the other two ellipticals were severely affected by the observing conditions, making their surface brightness profiles impossible to model. These galaxies were observed with between dithered and rotated exposures, each with an exposure time of 960 seconds, under bright Moon conditions, with typical lunar illuminations of . The sample analysed in this work consists of relatively isolated elliptical galaxies at , with stellar masses from the Nasa Sloan Atlas (NSA) 111Available from http://www.nsatlas.org.
The data reduction was carried out using the ESO MUSE pipeline (v2.6, Weilbacher et al., 2020) in the ESO Recipe Execution Tool (EsoRex) environment (ESO CPL Development Team, 2015). The master bias, flat field and wavelength calibrations for each CCD were created from the associated raw calibrations, and were applied to the raw science and standard-star observations as part of the pre-processing steps. Flux calibration was carried out using the standard star observations from the same nights as the science data, with the sky continuum measured directly from the science exposures with subsequent background subtraction. The individual exposures were aligned to a common frame and stacked to produce the final datacube. To alleviate residual sky contamination, we used the Zurich Atmosphere Purge code (ZAP, Soto et al., 2016). The coordinates and redshifts of the galaxies used in this analysis are listed in Table 1. For readability, these galaxies will henceforth be denoted as J020536 (MRK 1172), J205050, and J225546.
@@ 3 Methods
In this Section, we outline our methodology for decomposing the spectra of the elliptical galaxies using the BUlge-Disc Decomposition of IFU datacubes (BUDDI) software, followed by fitting the resulting spectra with pPXF. Additionally, we detail the construction of 2D stellar population maps utilising the Voronoi tessellation technique with Vorbin.
@@ 3.1 Galaxy decomposition with BUDDI
The light profiles of the elliptical galaxies in this sample were modelled using BUDDI (Johnston et al., 2017), which decomposes galaxies in IFU datacubes into multiple components and extracts each spectrum separately. This code functions as an IDL wrapper for GalfitM (Vika et al., 2013; Häußler et al., 2013; Vika et al., 2014; Häußler et al., 2022), extending its capabilities to handle the vast spectral dimension of IFU datacubes. While it would be ideal to freely model each image of a datacube with slices, this endeavour would require immense amounts of computational time and memory, that doubles with every additional component that is included in the model. BUDDI works around this limitation by first establishing the structure of the galaxy in a series of steps, then fitting only the magnitudes at each wavelength slice to construct the component spectra. This methodology enables clean separation of component spectra that would otherwise be contaminated by overlapping light from different galaxy components and, in some cases, the sky background (as demonstrated later in this Section).
@@
The PSF datacube, also log-rebinned in wavelength. In all three datacubes, postage stamps of stars in the field of view were created and stacked at each wavelength. A Gaussian profile was used to model them, constructing the PSF datacube. The full-width half maxima (FWHM) of the PSFs in our sample range from at the central band wavelength of 6166 Å.
The bad pixel mask datacube, that identifies and masks the pixels with no valid flux values. This also includes the masking of any bright sources in the field of view which can affect the fitting of the light profile of the target galaxy.
@@
Fitting the median “broad-band” image.
The datacube was collapsed in wavelength to a median-stacked white-light image. Within the BUDDI routine, this single image was fit using GalfitM with a single Sérsic component at first, and subsequently with two Sérsic components. The sky background amplitude and the gradient in and directions were allowed to be freely fit simultaneously - this was deemed necessary due to the inconducive observing conditions which had created a non-uniform sky background. -
@@
Fitting the “narrow-band” images.
This step fine-tunes the Chebyshev polynomials in GalfitM, which model the dependence of structural parameters (, , , PA) on wavelength for the galaxy in the single Sérsic model, and for both components in the double Sérsic model. The datacube was binned along wavelength into 10 “narrow-band” images (Johnston et al., 2022a, b; Jegatheesan et al., 2024) and fit simultaneously twice: first with complete freedom for the parameters (10th-order Chebyshev polynomial), and then manually choosing the appropriate polynomial order after visual inspection of the intermediate plots that BUDDI creates with the estimated structural parameters as a function of the wavelength. For each galaxy and component, the chosen polynomial orders for the parameters are listed below in Table 2: a first-order polynomial constrains the parameters to remain constant with wavelength, while a second-order polynomial allows for linear variation in wavelength. The magnitudes were allowed to have complete freedom during the fit, and the initial parameters were set to those measured in each preceding step.Table 2: Chebyshev polynomials introduced in the step fitting the narrow-band images of the datacube. Galaxy Comp PA J020536 single 2 2 1 1 inner 1 2 1 1 outer 1 2 1 1 J205050 single 2 2 1 1 inner 1 1 1 1 outer 1 1 1 1 J225546 single 1 2 1 1 inner 1 1 1 1 outer 1 2 1 1 -
@@
The final step makes use of the refined parameters from the previous narrow-band fits and keeps them fixed for each image slice while allowing only the magnitudes to vary in the fit. This estimates the magnitude, and therefore the flux of each component, in each image slice of the datacube, ultimately extracting their clean 1D spectra.

### v2

@@
Elliptical galaxies often exhibit complex assembly histories, and are presumed to typically form through a combination of rapid, early star formation and the subsequent accretion of material, often resulting from mergers with other galaxies. To investigate theories of spheroidal galaxy formation, the objective of this work is to analyse the star formation histories (SFHs) of a sample of three isolated elliptical galaxies in the local Universe observed with MUSE at . With BUDDI, we decompose the integral field unit (IFU) datacubes into two components with Sérsic profiles, which roughly correspond to the two phases of in situ and ex situ star formation. To constrain the mode of growth in these galaxies, we derived the mass and light-weighted stellar ages and metallicities, and created 2D stellar population maps of each component using pPXF. We reconstructed the mass and light-weighted SFHs to constrain the contribution of different stellar populations to the mass and luminosity of the components through cosmic time. Our results show that the ellipticals in this sample have experienced an early and rapid phase of star formation, either through a rapid dissipative collapse or gas-rich major mergers concentrated in the inner component, which contributes to % of the galaxy stellar mass. The co-dominant outer component, however, had assembled the bulk of its stellar mass shortly after the inner component did, through accretion via dry mergers and possible gas accretion. This premise is supported by our observations of the inner component being primarily composed of old and metal-rich stars. The outer component has a combination of old and intermediate-age stars, with a moderate spread in metallicities. These results are analysed through the lens of the two-phase scenario, a framework developed over the years to explain the formation histories of elliptical galaxies.
@@
Elliptical galaxies are early-type galaxies (ETGs) that are characterised by a spheroidal appearance and a seemingly featureless and plain morphology that lacks any intricate structure. In the local Universe, they are thought to be simple and massive entities that are ‘red and dead’, hosting old and quiescent stellar populations and meagre amounts of gas to support star formation. However, in reality, this morphology encompasses a wide range of stellar masses and physical sizes, from compact and dwarf ellipticals to the most massive giant ellipticals. Each type has distinct structural properties and surface brightness profiles, providing valuable insights into galaxy formation and evolution. The formation scenarios for elliptical galaxies highlight the complex interplay of physical processes, from rapid monolithic collapse to gradual hierarchical merging and accretion. For instance, an often-cited mechanism for the formation of massive ellipticals is gas-rich mergers between spiral galaxies (Toomre, 1977; Hopkins et al., 2008). These processes essentially shape the structure of the elliptical galaxies, and result in different stellar populations, stellar orbits, and kinematics. Modelling the surface brightness profiles of ellipticals is therefore a crucial step in uncovering their origins and evolutionary history (MacArthur et al., 2008; Vulcani et al., 2014; Spavone et al., 2017, 2021).
@@
Elliptical galaxies were traditionally modelled with a de Vaucouleurs profile (de Vaucouleurs, 1948), which is an law. This implies a galaxy with a central concentration that becomes more extended radially outward. While this captures the light profile of many ellipticals, improved instrumentation and spatial resolution over the years have revealed a large diversity of elliptical galaxies deviating from this law (Kormendy, 1977; King, 1978). The Sérsic profile (an law) was later introduced (Sersic, 1968), which modelled the surface brightness of the ellipticals more effectively, especially the inner regions of the galaxies. This was possible due to the freedom allowed for the Sérsic index, , which reverts to the de Vaucouleurs profile at . This flexibility that is incorporated into the Sérsic law is crucial because although ellipticals are indeed known for their high central concentration, the degree of concentration can vary from galaxy to galaxy. However, even the Sérsic profile might not fully represent the complex substructure that ellipticals are now expected to exhibit (Huang et al., 2013a, b; Lacerna et al., 2016), challenging the simple structure that is often assumed.
@@
Deep imaging has allowed for the observation of internal substructures in elliptical galaxies, including low-surface-brightness tidal features (van Dokkum, 2005; Tal et al., 2009; Lazar et al., 2023) stemming from their outskirts in cluster environments, or shells (Malin & Carter, 1980; Athanassoula & Bosma, 1985; Colbert et al., 2001; Bílek et al., 2016) around the galaxies. Substructures in the central regions of ellipticals have also been observed at different redshifts, such as embedded nuclear discs (Graham et al., 2016; Dullo et al., 2018; Lacerna et al., 2020; Deeley et al., 2023). Furthermore, ellipticals have been systematically observed to fall under one of two classes based on the steepness of their central surface brightness profiles: ‘core ellipticals’ with nearly flat central light profiles and ‘cuspy ellipticals’ with sharply peaked central light profiles. The core and cusp regions are considered to be photometrically distinct components of ellipticals (Huang et al., 2013a). This central component therefore deviates from the standard Sérsic law that attempts to describe the surface brightness profile of the entire elliptical galaxy with a single model, without accounting for the inner component. It is therefore important to decompose the components of deceptively simple elliptical galaxies to discern their physical relevance, their formation histories, and stellar populations. For instance, Huang et al. (2013a) found that their sample of 94 nearby elliptical galaxies in the Carnegie-Irvine Galaxy Survey (CGS) required three or even four components to optimise the modelling of the surface brightness profile. Apart from the structural analysis, a complementary study in Huang et al. (2013b) revealed the underlying physical implications of modelling multiple components. Similarly, Lacerna et al. (2016) found that their sample of 89 local elliptical galaxies at was best modelled with two to four components. More recently, Lima Neto et al. (2020) have demonstrated that two or three components are required to model the ellipticals in the NGC 4104 fossil group. The mechanisms leading to the formation of each component are more complex than can be explained by one of the two simple scenarios involving dissipational processes, as was previously mentioned.
@@
The two major pillars of elliptical galaxy formation scenarios are rapid dissipative collapse and hierarchical clustering. Over the decades, both numerical simulations and observations have been able to satisfactorily explain the structure of elliptical galaxies based on these theories. In the monolithic collapse scenario (Larson, 1974; Chiosi & Carraro, 2002), an elliptical galaxy is formed by the rapid gravitational collapse of a primordial gas cloud, followed by a substantial dissipation of energy. This theory predicts that ellipticals emerge at high redshifts () directly from the gas, with in situ stellar mass build-up, rather than from the accretion of stars formed ex situ (Ogando et al., 2005). The rapid monolithic collapse scenario has been corroborated by strong negative metallicity gradients observed in elliptical galaxies (Domínguez Sánchez et al., 2019; Parikh et al., 2019; Goddard et al., 2017; Riffel et al., 2023), owing to the fact that the metals created during supernova explosions are retained by a strong-enough gravitational potential that pushes the metal-enhanced gas towards the centre of the galaxy, where the new stars form at a higher metallicity compared to the outer regions of the galaxy. The hierarchical clustering scenario (Toomre & Toomre, 1972; Cole et al., 2000; Hopkins, 2009; Avila-Reese et al., 2014), on the other hand, predicts that the present-day galaxies formed through major and minor mergers and accretion of galaxies over time. This theory has also been supported by simulations and observations, where a flat or shallow metallicity gradient was exhibited by ellipticals (González Delgado et al., 2015; Taylor & Kobayashi, 2017; Benedetti et al., 2023), instead of the steep negative gradient expected from the monolithic collapse. Pure merger events are presumed to obliterate traces of the previous stellar history of the galaxy, including metallicity gradients.
@@
To better explain the mixed outcomes of elliptical galaxy observations, a hybrid scenario was proposed (Kormendy, 1989), according to which both mechanisms play a role in the formation of these remarkably complex galaxies. In the last decades, a two-phase scenario has been gaining traction, challenging the classical pictures of elliptical galaxy formation from either mergers or dissipative gas collapse (Kormendy, 1989). The two-phase scenario (Oser et al., 2010; Johansson et al., 2012; Huang et al., 2013a, b) suggests that the first phase of dissipational processes dominate at high redshift through cold accretion and wet mergers of gas-rich galaxies, and that later, at lower redshifts, the galactic evolution is driven by delayed non-dissipational dry mergers. The presence of ‘red nuggets’ at high redshifts, , detected in several studies (Cimatti et al., 2004; Daddi et al., 2005; Trujillo et al., 2007; Toft et al., 2007; van Dokkum et al., 2008; Damjanov et al., 2009) already indicates an early onset of rapid star formation that settles into a compact core. Moreover, a smaller population has been identified of red nuggets that have survived to the present day, having evolved in isolation without perturbations or interactions (Ferré-Mateu et al., 2017; Siudek et al., 2023; Moura et al., 2024; Spiniello et al., 2024). Most of these red nuggets eventually accumulate more stellar mass by accreting and merging with smaller systems to acquire the structure that is observed in present-day elliptical galaxies (Miller & Smith, 1980; Naab et al., 2006; Oser et al., 2010, 2012), while some ellipticals seem to have preserved their compact high- progenitor (Barbosa et al., 2021).
@@
Given the intricate structure of elliptical galaxies as both observations and numerical simulations clearly support, we delve into the multi-component context of these galaxies in this paper. We make use of integral field spectroscopy (IFS) with the Multi Unit Spectroscopic Explorer (MUSE) observations of three elliptical galaxies in the local Universe, and perform 2D spectro-photometric decomposition to retrieve their structural parameters, as well as to extract their spectra and stellar populations to interpret their physical significance. The decomposition of a sample of three galaxies serves as a pilot study for further exploring the diverse properties that are inherent to elliptical galaxies over a range of stellar masses, redshifts, and environments.
@@
This paper is organised as follows. Section 2 outlines the data and instrument used in the IFS observations for this study. Section 3 describes the methods of galaxy decomposition with BUDDI, full spectral fitting with pPXF, and Voronoi binning with Vorbin. The results of this work are highlighted in Sect. 4 in the context of structural parameters, stellar population analysis, and the star formation histories (SFHs) of each component. Section 5 provides a discussion on the physical implications and future directions of this study, and we draw our conclusions in Sect. 6. Throughout this work, we adopt a flat CDM cosmology with (Planck Collaboration et al., 2016).
@@ 2 Data and observations
MUSE (Bacon et al., 2010) is an optical integral-field spectrograph mounted at the 8.2m ESO Very Large Telescope (VLT) in Chile. It covers the nominal wavelength range of Å with spectral sampling of 1.25 Å and a field of view, and spatial resolution of 0.2″/pixel in the Wide Field Mode (WFM). The MUSE spectrograph has a spectral resolving power ranging from R1770 at 4800 Å to R3590 at 9300 Å. The three ellipticals were observed under Program ID 099.B-0411 (PI. Johnston) between July and October 2017 using the wide field mode with no adaptive optics (WFM-NoAO-N). The targets initially proposed for observation consisted of 48 galaxies selected from SDSS DR7, representing all major Hubble types from Nair & Abraham (2010). The selected targets spanned a range of Sérsic indices and B/T ratios, with an inclination derived from Simard et al. (2011), which allowed us to clearly discern the components of the galaxies. Moreover, the sample was chosen to only include unbarred galaxies, with no signatures of recent interactions, such that each galaxy can be cleanly decomposed into its major components. However, being a filler programme observed under sub-optimal weather conditions, only a small set of five elliptical galaxies with a higher set priority were observed. From this, we consider only three galaxies to be suitable for analysis in this work, as the other two ellipticals were severely affected by the observing conditions, making their surface brightness profiles impossible to model. These galaxies were observed with between dithered and rotated exposures, each with an exposure time of 960 seconds, under bright moon conditions, with typical lunar illuminations of . The sample analysed in this work consists of relatively isolated elliptical galaxies at , with stellar masses of from the Nasa Sloan Atlas (NSA) 111Available from http://www.nsatlas.org.
The data reduction was carried out using the ESO MUSE pipeline (v2.6, Weilbacher et al., 2020) in the ESO Recipe Execution Tool (EsoRex) environment (ESO CPL Development Team, 2015). The master bias, flat field, and wavelength calibrations for each CCD were created from the associated raw calibrations, and were applied to the raw science and standard-star observations as part of the pre-processing steps. Flux calibration was carried out using the standard star observations from the same nights as the science data, with the sky continuum measured directly from the science exposures with subsequent background subtraction. The individual exposures were aligned to a common frame and stacked to produce the final datacube. To alleviate residual sky contamination, we used the Zurich Atmosphere Purge code (ZAP, Soto et al., 2016). The co-ordinates and redshifts of the galaxies used in this analysis are listed in Table 1. For readability, these galaxies will henceforth be denoted as J020536 (MRK 1172), J205050, and J225546.
@@ 3 Methods
In this section, we outline our methodology for decomposing the spectra of the elliptical galaxies using the BUlge-Disc Decomposition of integral field unit (IFU) datacubes (BUDDI) software, followed by fitting the resulting spectra with pPXF. Additionally, we detail the construction of 2D stellar population maps utilising the Voronoi tessellation technique with Vorbin.
@@ 3.1 Galaxy decomposition with BUDDI
The light profiles of the elliptical galaxies in this sample were modelled using BUDDI (Johnston et al., 2017), which decomposes galaxies in IFU datacubes into multiple components and extracts each spectrum separately. This code functions as an IDL wrapper for GalfitM (Vika et al., 2013; Häußler et al., 2013; Vika et al., 2014; Häußler et al., 2022), extending its capabilities to handle the vast spectral dimension of IFU datacubes. While it would be ideal to freely model each image of a datacube with slices, this endeavour would require immense amounts of computational time and memory, which doubles with every additional component that is included in the model. BUDDI works around this limitation by first establishing the structure of the galaxy in a series of steps, then fitting only the magnitudes at each wavelength slice to construct the component spectra. This methodology enables clean separation of component spectra that would otherwise be contaminated by overlapping light from different galaxy components and, in some cases, the sky background (as is demonstrated later in this section).
@@
The PSF datacube, which has also been log-rebinned in wavelength. In all three datacubes, postage stamps of stars in the field of view were created and stacked at each wavelength. A Gaussian profile was used to model them, constructing the PSF datacube. The full widths at half maximum of the PSFs in our sample range from at the central band wavelength of 6166 Å.
The bad pixel mask datacube, which identifies and masks the pixels with no valid flux values. This also includes the masking of any bright sources in the field of view that can affect the fitting of the light profile of the target galaxy.
@@
Fitting the median ‘broad-band’ image.
The datacube was collapsed in wavelength to a median-stacked white-light image. Within the BUDDI routine, this single image was fitted using GalfitM with a single Sérsic component at first, and subsequently with two Sérsic components. The sky background amplitude and the gradient in the and directions were allowed to be freely fitted simultaneously - this was deemed necessary due to the inconducive observing conditions, which had created a non-uniform sky background. -
@@
Fitting the ‘narrow-band’ images.
This step fine-tunes the Chebyshev polynomials in GalfitM, which model the dependence of structural parameters (, , , PA) on wavelength for the galaxy in the single Sérsic model, and for both components in the double Sérsic model. The datacube was binned along wavelength into 10 narrow-band images (Johnston et al., 2022a, b; Jegatheesan et al., 2024) and fitted simultaneously twice: first with complete freedom for the parameters (10th-order Chebyshev polynomial), and then manually choosing the appropriate polynomial order after visual inspection of the intermediate plots that BUDDI creates with the estimated structural parameters as a function of the wavelength. For each galaxy and component, the chosen polynomial orders for the parameters are listed below in Table 2: a first-order polynomial constrains the parameters to remain constant with wavelength, while a second-order polynomial allows for linear variation in wavelength. The magnitudes were allowed to have complete freedom during the fit, and the initial parameters were set to those measured in each preceding step.Table 2: Chebyshev polynomials introduced in the step fitting the narrow-band images of the datacube. Galaxy Comp PA J020536 single 2 2 1 1 inner 1 2 1 1 outer 1 2 1 1 J205050 single 2 2 1 1 inner 1 1 1 1 outer 1 1 1 1 J225546 single 1 2 1 1 inner 1 1 1 1 outer 1 2 1 1 -
@@
The final step makes use of the refined parameters from the previous narrow-band fits and keeps them fixed for each image slice, while allowing only the magnitudes to vary in the fit. This estimates the magnitude, and therefore the flux of each component, in each image slice of the datacube, ultimately extracting their clean 1D spectra.